1-

In a particular lake, all fish weights have a mean of μ = 17.9 g. What is the probability that a sample of n 11 fish caughtntis lake will have a mean weight between 17.1 g and 18.1 g? (Assume that s 2.3 g

2-
Answer #1

1)

We would be using the below formula: -

z=x-\mu/\sigma/\sqrtn

z=17.1-17.9/2.3/\sqrt11

z=-1.153

z=x-\mu/\sigma

z=18.1-17.9/2.3/\sqrt11

z=0.2884

Looking at the z table, we get the values as 0.1251 for -1.153 and 0.6141 for 0.2884. Hence, 0.6141-0.1251=0.489 is the final answer